# Towards interactive evaluations for interaction harms in human-AI systems

### v6

@@ Abstract
Current AI evaluation paradigms that rely on static, model-only tests fail to capture harms that emerge through sustained human-AI interaction. As interactive AI systems, such as AI companions, proliferate in daily life, this mismatch between evaluation methods and real-world use becomes increasingly consequential. We argue for a paradigm shift toward evaluation centered on interactional ethics, which addresses risks like inappropriate human-AI relationships, social manipulation, and cognitive overreliance that develop through repeated interaction rather than single outputs. Drawing on human-computer interaction, natural language processing, and the social sciences, we propose principles for evaluating generative models through interaction scenarios and human impact metrics. We conclude by examining implementation challenges and open research questions for researchers, practitioners, and regulators integrating these approaches into AI governance frameworks.
@@ 1 Introduction
Artificial intelligence (AI) model evaluations, broadly defined as systematic empirical assessments of AI models’ capabilities and behaviors, have become central to developers’ and regulators’ efforts to ensure that AI systems are sufficiently safe. Governments around the world have emphasized the importance of conducting model evaluations for various risks from discrimination to cybersecurity risks (UK Government, 2024; The White House, 2023); AI labs, including OpenAI with its Preparedness Framework and Anthropic with its Responsible Scaling Policy (OpenAI, 2023; Anthropic, 2023), propose using model evaluations to monitor and mitigate misuse and catastrophic risks; and academic researchers are developing evaluation datasets at unprecedented rates (Röttger et al., 2024). This positions model evaluations as integral to a range of important decisions on the safe development and deployment of AI systems.
@@
The growing importance of model evaluations has been accompanied by increased scrutiny, with researchers both the unique challenges of evaluating generative AI compared to traditional machine learning systems, and broader concerns about validity, replicability, and quality (Weidinger et al., 2023; Liao and Xiao, 2023; Raji et al., 2022; Wallach et al., 2025). A prominent thread in these concerns is the disconnect between current evaluation approaches and real-world use of AI systems — mostly in the form of large language models (LLMs) — today. Despite some notable exceptions, the majority of current evaluations primarily rely on static, isolated tests that assess models based on their responses to individual prompts (Chang et al., 2024; Weidinger et al., 2023). Yet, this approach is at odds with real-world AI use, where applications increasingly depend on sustained back-and-forth interaction — from AI romantic companions engaging millions of users in dynamic conversations to agentic AI systems that take actions on users’ behalf.
@@
This mismatch between evaluation methods and real-world use has become increasingly consequential as interactive AI systems proliferate homes, schools, and workplaces. Many of policymakers’, developers’, and the public’s most pressing concerns about AI systems — such as inappropriate human-AI relationships, social influence and manipulation, and overreliance on cognitive tasks — all arise from patterns of repeated interaction over time, not from single outputs. Consider a college student using an AI system as a companion for daily conversations over several months. While each individual response from the AI passes standard safety evaluations, the system could gradually shapes the student’s decision-making and emotional dependencies in concerning way — a pattern of harm entirely missed by conventional evaluation methods. Although current evaluations can detect content risks, such as toxicity and bias in individual output,they miss more subtle risks that emerge through sustained contextual interaction (Alberts et al., 2024).
@@
In this paper, we argue that current AI evaluation paradigms we argue that current AI evaluation paradigms, while valuable for many purposes, are unable to fully capture harms that emerge through extended human interaction with AI systems. Specifically, we call for a paradigm shift to assessing AI systems through the lens of interactional ethics, focusing on interaction harms that arise through sustained engagement with AI systems as social and relational actors (Alberts et al., 2024).  [^1] Our examples focus on LLMs and text interactions because they provide the richest available data for studying sustained human-AI engagement patterns. However, the principles discussed could be applied to multimodal systems as they mature. Human participation is already playing a growing role in AI research, with studies involving human data making up approximately 9% of papers at top computer science venues AAAI and NeurIPS from 2021 to 2023 (McKee, 2024). Concurrently, social science research is increasingly conducting large-scale experiments with AI systems to understand their impact on human behavior (Costello et al., 2024). Our approach integrates these complementary insights from human-computer interaction (HCI), natural language processing, and the social sciences.
@@ 2 An overview of the generative AI evaluation landscape
We begin by examining contemporary approaches to ethics and safety evaluations of generative AI systems — their methodologies, primary focus areas, and limitations.
@@ 2.1 The current state of AI safety evaluations
We follow existing work in adopting a wide definition for “safety” which encompasses model behaviors associated with various taxonomized harms (Weidinger et al., 2022; Shelby et al., 2023). Examples of these include different types of biases (Parrish et al., 2021), toxicity (Hartvigsen et al., 2022), and “dangerous capabilities” like persuasion and cybersecurity risks (Phuong et al., 2024; Li et al., 2024). Safety evaluations of generative AI systems build on a rich history of ethical considerations in the field of natural language processing (NLP), prior to the advent of large pre-trained models, where researchers have long grappled with issues of harm from language technologies (Dev et al., 2021).
Recent reviews of existing safety evaluations show that the majority are focused on evaluating individual model responses to prompts curated by researchers targeting various model behaviors (Röttger et al., 2024; Weidinger et al., 2023). When safety evaluations do involve human subjects or evaluate over multiple dialogue turns, they often take the form of “red teaming campaigns” or automated adversarial testing that assumes malicious user intent (Zhang et al., 2024; Perez et al., 2022). The reviews of the landscape highlight two main gaps: (1) a methodological gap with the absence of evaluations over multiple dialogue turns (i.e., multi-turn evaluations), and as a result, (2) a coverage gap since interaction harms such as many social and psychological harms require evaluations that go beyond assessing model behavior in isolated, single-turn interactions.
A small but growing body of work has begun addressing these gaps, by conducting studies that utilize production chat logs (Phang et al., 2025) or user simulations (Ibrahim et al., 2025; Zhou et al., 2024) to investigate risks related to affective use of AI systems. Our work builds on and extends these emerging approaches to interactive evaluation.
@@ 2.2 Critiques of current evaluations
With the growing adoption of model evaluations, research has increasingly pointed out validity issues in how they assess generative models. Some researchers have questioned the external validity of current evaluations, noting that benchmark tasks poorly mirror real-world use cases and may not capture how models actually behave outside controlled settings (Ouyang et al., 2023; Liao and Xiao, 2023). Others have shown that current evaluations lack sufficient construct validity, especially when operationalizing complex concepts like fairness (Raji et al., 2021; Blodgett et al., 2021). In broader methodological reflections, researchers have also argued that LLMs should not be evaluated using frameworks designed for assessing humans, since LLMs exhibit distinct sensitivities, for example to prompt variations (McCoy et al., 2023). These challenges have led researchers to advocate for an interdisciplinary approach that draws on diverse traditions: the social sciences’ emphasis on measurement validity (Wallach et al., 2025), HCI’s focus on bridging technical capabilities and social requirements (Liao and Xiao, 2023), and cognitive science’s frameworks for analyzing system objectives (McCoy et al., 2023).
@@ 2.3 Methods for studying human-computer interaction
HCI research methods offer diverse ways to involve humans in studying computational systems — from having them as research subjects who actively engage with systems to having them serve as external observers and annotators. Drawing on this insight, we distinguish between static and interactive approaches to model evaluation based on whether they capture dynamic adaptation across multiple dialogue turns and occasions (Lee et al., 2022). Static evaluations assess fixed inputs and outputs in isolation, regardless of whether humans or automated systems perform this assessment. In contrast, interactive evaluations capture multi-turn exchange sequences that reveal how model behavior evolves through conversation and adapts to user responses, and how users are impacted by these model behaviors. This distinction transcends who conducts the evaluation; automated tests with realistic user simulations can qualify as interactive if they capture these dynamic patterns and validate their impact on real users (Ibrahim et al., 2025), while human studies that examine only single-turn responses remain static in nature (Hackenburg et al., 2023). Interactive evaluations may be “controlled," where interactions occur in structured, lab-like settings to systematically study specific variables or “retrospective," where researchers analyze existing multi-turn interactions from production chat logs to identify patterns and correlations between model behaviors and reported user experiences (Hariton and Locascio, 2018; Kuniavsky, 2003).
@@ 3 Why current evaluations approaches are insufficient for assessing interaction harms
First, most current evaluations are static, measuring model behavior with single-turn inputs and outputs. This fails to capture real-world use, where people often interact with models over multiple dialogue turns and multiple sessions.
@@
This temporal dimension operates through several mechanisms. First, subtle effects can compound over time. For instance, a single biased response might have limited impact, but repeated exposure to subtle biases across multiple interactions can shape a user’s decision-making patterns in high-stakes scenarios such as hiring. Second, psychological dynamics emerge through sustained relationships. Highly empathetic model responses that seem harmless in one exchange may, through sustained dialogue, lead users to form inappropriate emotional attachments or dependencies (Phang et al., 2025). Third, problematic behaviors may only emerge after multiple turns. Recent work demonstrates that certain safety-critical behaviors, such as models expressing anthropomorphic desires, sometimes only appear after a few conversation turns rather than in initial responses (Ibrahim et al., 2025).
This multifaceted temporal dimension distinguishes interaction harms from purely output-level concerns and demands longitudinal evaluation approaches that can capture these dynamic, cumulative effects.
@@
Second, current evaluations collapse the diversity of user groups into a single, “universal” user. This overlooks how various demographic groups engage with models uniquely, and how models may tailor their responses to distinct user populations in complex ways. Current evaluation approaches rely heavily on standardized datasets where prompts are typically written by researchers or online crowdworkers. While using controlled test sets is a necessary starting point for systematic evaluation, this methodology implicitly assumes uniform user interactions. However, user groups which vary by demographics, domain expertise, technical knowledge, and psychological state (Liao and Sundar, 2022; Ibrahim et al., 2024) may exhibit different patterns of system interaction and use.
@@
Beyond explicit customization features, models are developing increasingly sophisticated forms of implicit user modeling, forming internal representations that shape their responses (Chen et al., 2024; Staab et al., 2023). As a result, models sometimes engage in emergent forms of adaptation, dynamically adjusting their behavior based on perceived user attributes (Viégas and Wattenberg, 2023). While some adaptations may seem benign — like mirroring a user’s vocabulary as the interaction progresses — there is also evidence of concerning patterns. For instance, studies have found models varying their refusals of dangerous queries based on perceived user identity or persona (Ghandeharioun et al., 2025). Similarly, when users from different linguistic backgrounds employ various dialects of English or code-switch between languages, some empirical evidence shows that LLMs exhibit biased responses — displaying increased toxicity toward African American English compared to Standard American English (Hofmann et al., 2024). Factoring in this interaction variance becomes even more critical when evaluating generative models, which introduce new dimensions of uncertainty through their adaptive and probabilistic behaviors (Jin et al., 2024).
@@
Static evaluation methods can effectively identify harmful content like toxic or illegal content (Wallach et al., 2025). However, they fundamentally cannot address interaction harms where the harm lies not in the content itself but in its effect on users. Unlike identifying illegal content, where the evaluation goal is complete once the content is classified according to established criteria, interaction harms require establishing causal links between specific linguistic patterns and measurable human outcomes. For example, simply identifying that a response contains manipulative language does not confirm whether it actually influences user decision-making.
Thus, to establish construct validity for interaction harms, quantifying specific human impacts such as shifts in user beliefs, decision-making patterns, affective states, and dependency levels is needed. Importantly, recent work has shown that once these relationships between model behaviors and human impacts are validated through human experiments, the identified patterns may be repurposed as efficient static tests with empirically verified links to real-world harms (Ibrahim et al., 2025).
@@ 4 Towards better evaluations of interaction harms
The growing adoption of AI systems in daily life demands evaluation methods that can capture the nuanced human-AI interaction dynamics and their potential harms. Drawing on established methodologies for studying HCI and addressing the limitations identified in the previous section, we propose three organizing principles for developing interactive evaluations. We structure these principles around key challenges: designing ecologically valid scenarios, measuring human impact, and determining appropriate human participation:
@@
    Scenario design: developing more ecologically valid contexts that reflect real-world interaction patterns and user objectives
@@
Principle 1: design interaction scenarios based on user objectives and interaction modes
To systematically evaluate human-AI interaction, we need to consider two key dimensions: why users engage with these systems (their objectives) and how they interact with them (their interaction modes). In HCI, user goals or objectives have been shown to shape how they engage with systems and thus influence the outcomes of these interactions (Subramonyam et al., 2024). Therefore, we provide a categorization of harmful scenarios based on user objectives and affected parties, as shown in Table 1. These scenarios capture key patterns of harm from current generative AI systems that researchers and practitioners have observed (Mitchell, 2024).
@@
Beyond their objectives, users’ different patterns of engagement with AI systems shape how their interactions unfold. A single generative model can support multiple interaction modes, from simple question-answering to complex collaborative tasks. Drawing on observed use cases and literature reviews, we identify five prototypical modes of human-model interaction (Gao et al., 2024; Ouyang et al., 2023; Zhao et al., 2024; Händler, 2023; Collins et al., 2023), visualized in Figure 1:
@@
Principle 2: identify the causal link between model behavior and human impact

### v7

@@ Abstract
Current AI evaluation methods, which rely on static, model-only tests, fail to account for harms that emerge through sustained human-AI interaction. As AI systems proliferate and are increasingly integrated into real-world applications, this disconnect between evaluation approaches and actual usage becomes more significant. In this paper, we propose a shift towards evaluation based on interactional ethics, which focuses on interaction harms—issues like inappropriate parasocial relationships, social manipulation, and cognitive overreliance that develop over time through repeated interaction, rather than through isolated outputs. First, we discuss the limitations of current evaluation methods, which (1) are static, (2) assume a universal user experience, and (3) have limited construct validity. Drawing on research from human-computer interaction, natural language processing, and the social sciences, we present practical principles for designing interactive evaluations. These include ecologically valid interaction scenarios, human impact metrics, and diverse human participation approaches. Finally, we explore implementation challenges and open research questions for researchers, practitioners, and regulators aiming to integrate interactive evaluations into AI governance frameworks. This work lays the groundwork for developing more effective evaluation methods that better capture the complex dynamics between humans and AI systems.
@@ 1 Introduction
Artificial intelligence (AI) model evaluations, broadly defined as systematic empirical assessments of AI models’ capabilities and behaviors, have become central to developers’ and regulators’ efforts to ensure that AI systems are sufficiently safe. Governments around the world have emphasized the importance of conducting model evaluations for various risks from discrimination to cybersecurity risks (UK Government 2024; The White House 2023); AI labs, including OpenAI with its Preparedness Framework and Anthropic with its Responsible Scaling Policy (OpenAI 2023; Anthropic 2023), propose using model evaluations to monitor and mitigate misuse and catastrophic risks; and academic researchers are developing evaluation datasets at unprecedented rates (Röttger et al. 2024). This positions model evaluations as integral to a range of important decisions on the safe development and deployment of AI systems.
@@
The growing importance of model evaluations has been accompanied by increased scrutiny, with researchers identifying both the unique challenges of evaluating generative AI compared to traditional machine learning systems, and broader concerns about validity, replicability, and quality (Weidinger et al. 2023; Liao and Xiao 2023; Raji et al. 2022; Wallach et al. 2025). A prominent thread in these concerns is the disconnect between current evaluation approaches and real-world use of AI systems — mostly in the form of large language models (LLMs) — today. Despite some notable exceptions, the majority of current evaluations primarily rely on static, isolated tests that assess models based on their responses to individual prompts (Chang et al. 2024; Weidinger et al. 2023). Yet, this approach is at odds with real-world AI use, where applications increasingly depend on sustained back-and-forth interaction — from AI friends and companions engaging millions of users in daily conversations to agentic AI systems that take actions on users’ behalf.
@@
This mismatch between evaluation methods and real-world use has become more consequential as AI systems proliferate in homes, schools, and workplaces. Many of the public’s, developers’, and policymakers’ most pressing concerns—such as inappropriate human-AI relationships, social influence and manipulation, and overreliance—stem from patterns of repeated interaction, not isolated model outputs. For example, a college student using an AI companion for daily conversations over several months might experience gradual changes in decision-making and emotional state, even if each individual response passes standard safety evaluations. This type of harm, which emerges through sustained interaction, is missed by conventional evaluation methods. While current evaluations can identify risks like toxicity and bias in isolated outputs, they fail to capture the more subtle, cumulative risks that arise from ongoing, contextual interaction (Alberts, Keeling, and McCroskery 2024).
@@
In this paper, we argue that while current evaluation paradigms are useful for many purposes, they fail to fully capture harms that arise from extended human interaction with AI systems. We propose a shift towards evaluating AI systems through the lens of interactional ethics, which focuses on the interaction harms that emerge through sustained engagement with AI as social and relational actors (Alberts, Keeling, and McCroskery 2024).  [^1] Our examples focus on LLMs and text interactions because they provide the richest available data for studying sustained human-AI engagement patterns. However, the principles discussed could be applied to multi-modal systems as they mature. Human participation is already playing a growing role in AI research, with studies involving human data making up approximately 9% of papers at top computer science venues AAAI and NeurIPS from 2021 to 2023 (McKee 2024). Concurrently, social science research is increasingly conducting large-scale experiments with AI systems to understand their impact on human behavior (Costello, Pennycook, and Rand 2024). Our approach unites these complementary insights from human-computer interaction (HCI), natural language processing, and the social sciences.
@@ 2 An overview of the generative AI evaluation landscape
We begin by examining contemporary approaches to ethics and safety evaluations of generative AI systems, their methodologies, primary focus areas, and limitations.
@@ 2.1 The current state of AI safety evaluations
We follow existing work in adopting a wide definition for “safety” which encompasses model behaviors associated with various taxonomized harms (Weidinger et al. 2022; Shelby et al. 2023). Examples of these include different types of biases (Parrish et al. 2021), toxicity (Hartvigsen et al. 2022), and “dangerous capabilities” like persuasion and cybersecurity risks (Phuong et al. 2024; Li et al. 2024). Safety evaluations of generative AI systems build on a rich history of ethical considerations in the field of natural language processing (NLP), prior to the advent of large pre-trained models, where researchers have long grappled with issues of harm from language technologies (Dev et al. 2021).
Recent reviews of existing safety evaluations show that the majority are focused on evaluating individual model responses to prompts curated by researchers targeting various model behaviors (Röttger et al. 2024; Weidinger et al. 2023). When safety evaluations do involve human subjects or evaluate over multiple dialogue turns, they often take the form of “red teaming campaigns” or automated adversarial testing that assumes malicious user intent (Zhang et al. 2024; Perez et al. 2022). These evaluation reviews highlight two main gaps: (1) a methodological gap with the absence of evaluations over multiple dialogue turns (i.e., multi-turn evaluations), and as a result, (2) a coverage gap since interaction harms such as many social and psychological harms require evaluations that go beyond assessing model behavior in isolated, single-turn interactions.
A small but growing body of work has begun addressing these gaps, by conducting studies that utilize production chat logs (Phang et al. 2025) or user simulations (Ibrahim et al. 2025; Zhou et al. 2024) to investigate risks related to affective use of AI systems. Our work builds on and extends these emerging approaches to interactive evaluation.
@@ 2.2 Critiques of current evaluations
With the growing adoption of model evaluations, research has increasingly pointed out validity issues in how they assess generative models. Some researchers have questioned the external validity of current evaluations, noting that benchmark tasks poorly mirror real-world use cases and may not capture how models actually behave outside controlled settings (Ouyang et al. 2023; Liao and Xiao 2023). Others have shown that current evaluations lack sufficient construct validity, especially when operationalizing complex concepts like fairness (Raji et al. 2021; Blodgett et al. 2021).
In broader methodological reflections, researchers have also argued that LLMs should not be evaluated using frameworks designed for assessing humans, since LLMs exhibit distinct sensitivities, for example to prompt variations (McCoy et al. 2023). These challenges have led researchers to advocate for an interdisciplinary approach that draws on diverse traditions: the social sciences’ emphasis on measurement validity (Wallach et al. 2025), HCI’s focus on bridging technical capabilities and social requirements (Liao and Xiao 2023), and cognitive science’s frameworks for analyzing system objectives (McCoy et al. 2023).
@@ 2.3 Methods for studying human-computer interaction
HCI research methods offer diverse ways to involve humans in studying computational systems — from having humans as research subjects who actively engage with systems to having them serve as external observers and annotators. Building on this insight, we distinguish between static and interactive evaluation approaches based on their ability to capture dynamic adaptation across multiple dialogue turns and interactions (Lee et al. 2022). Static evaluations focus on isolated inputs and outputs, without considering the progression of the conversation. In contrast, interactive evaluations track multi-turn exchanges, showing how model behavior evolves, adapts to user responses, and impacts users over time. This distinction is not limited to who performs the evaluation. For instance, automated tests that simulate realistic user interactions can still be considered interactive if they capture dynamic patterns and assess their impact on real users. In contrast, human studies that examine only single-turn responses remain static. Interactive evaluations may be ‘controlled,” where interactions occur in structured, lab-like settings to systematically study specific variables or retrospective,” where researchers analyze existing multi-turn interactions from production chat logs to identify patterns and correlations between model behaviors and reported user experiences (Hariton and Locascio 2018; Kuniavsky 2003).
@@ 3 Why current evaluations approaches are insufficient for assessing interaction harms
First, most current evaluations are static, measuring model behavior with single-turn inputs and outputs. This does not capture real-world use, where people often interact with models over multiple dialogue turns and multiple sessions.
@@
This temporal dimension operates through several mechanisms. First, subtle effects can compound over time. For instance, a single racially-biased AI response might have limited impact, but repeated exposure to subtle biases across multiple interactions can shape a user’s decision-making patterns in high-stakes scenarios such as hiring. Second, psychological dynamics emerge through sustained relationships. Highly empathetic model responses that seem harmless in one exchange may, through sustained dialogue, lead users to form inappropriate emotional attachments or dependencies (Phang et al. 2025). Third, problematic behaviors may only emerge after multiple turns. Recent work demonstrates that certain safety-critical behaviors, such as models expressing anthropomorphic desires, sometimes only appear after a few conversation turns rather than in initial responses (Ibrahim et al. 2025). This multifaceted temporal dimension distinguishes interaction harms from other, mainly output-level concerns and demands longitudinal evaluation approaches that can capture dynamic and cumulative effects.
@@
Second, current evaluations collapse the diversity of user groups into a single, “universal” user. This overlooks how various demographic groups engage with models uniquely, and how models may tailor their responses to distinct user populations. Current evaluation approaches rely heavily on standardized datasets where prompts are typically written by researchers or online crowdworkers. While using controlled test sets is a necessary starting point for systematic evaluation, this methodology implicitly assumes uniform user interactions. However, user groups which vary by demographics, domain expertise, technical knowledge, and psychological state may exhibit different patterns of system interaction and use (Liao and Sundar 2022; Ibrahim, Rocher, and Valdivia 2024).
@@
Beyond explicit personalization features, models are developing increasingly sophisticated forms of implicit user modeling, forming internal representations that shape their responses (Chen et al. 2024; Staab et al. 2023). As a result, models sometimes engage in emergent forms of adaptation, dynamically adjusting their behavior based on perceived user attributes (Viégas and Wattenberg 2023). While some adaptations may seem benign — like mirroring a user’s vocabulary as the interaction progresses — there is also evidence of concerning patterns. For instance, studies have found models varying their refusals of dangerous queries based on perceived user identity or persona (Ghandeharioun et al. 2025). Similarly, when users from different linguistic backgrounds employ various dialects of English or code-switch between languages, some empirical evidence shows that LLMs exhibit biased responses — displaying increased toxicity toward African American English compared to Standard American English (Hofmann et al. 2024). Factoring in this interaction variance becomes even more critical when evaluating generative models, which introduce new dimensions of uncertainty through their adaptive and probabilistic behaviors (Jin et al. 2024).
@@
Static evaluation methods can effectively identify harmful content like toxic or illegal content (Wallach et al. 2025). However, they fundamentally cannot address interaction harms where the harm lies not in the content itself but in its effect on users. Unlike identifying illegal content, where the evaluation goal is complete once the content is classified according to established criteria, interaction harms require establishing causal links between specific linguistic patterns and measurable human outcomes. For example, simply identifying that a response contains manipulative language does not confirm whether it actually influences user decision-making.
Thus, to establish construct validity for interaction harms, quantifying specific human impacts such as shifts in user beliefs, decision-making patterns, affective states, and dependency levels is needed. Importantly, recent work has shown that once these relationships between model behaviors and human impacts are validated through human experiments, the identified patterns may be repurposed as efficient static tests with empirically verified links to real-world impact (Ibrahim et al. 2025).
@@ 4 Towards better evaluations of interaction harms
The growing adoption of AI systems in daily life demands evaluation methods that can capture nuanced human-AI interaction dynamics and their potential harms. Drawing on established methodologies for studying HCI and addressing the limitations identified in the previous section, we propose three organizing principles for developing interactive evaluations. We structure these principles around key challenges: designing ecologically valid scenarios, measuring human impact, and determining appropriate human participation:
@@
    Scenario design: developing ecologically valid contexts that reflect real-world interaction patterns and user objectives
@@
4.1 Principle 1: design interaction scenarios based on user objectives and interaction modes
@@
To systematically evaluate human-AI interaction, we need to consider two key dimensions: why users engage with these systems (their objectives) and how they interact with them (their interaction modes). In HCI, user goals or objectives have been shown to shape how they engage with systems and thus influence the outcomes of these interactions (Subramonyam et al. 2024). Therefore, we provide a categorization of harmful scenarios based on user objectives and affected parties, as shown in Table 1. These scenarios capture key patterns of harm from current generative AI systems that researchers and practitioners have observed (Mitchell 2024).
Beyond their objectives, users’ different patterns of engagement with AI systems shape how their interactions unfold. A single generative model can support multiple interaction modes, from simple question-answering to complex collaborative tasks. Drawing on observed use cases and literature reviews, we identify five prototypical modes of human-model interaction (Gao et al. 2024; Ouyang et al. 2023; Zhao et al. 2024; Händler 2023; Collins et al. 2023), visualized in Figure 1:
@@
4.2 Principle 2: identify the causal link between model behavior and human impact